A carnot engine has its source at 120°C and its sink is maintained at a constant temperature by means of ice at 0°C. If it is working at the rate of 175 watts, how much ice will melt in two minutes?
1
Expert's answer
2020-08-21T10:59:20-0400

Given data

Temperature of source "(T"source)=1200C+273

=393 K

Temperature of sink (Tsink)=00C+273

=273 K

Working rate (power) =175 watts

Workdone by carnot engine in 2 minutes

"W=P\times t"

"W=175\times2 \times60"

"W=21000 J"

efficiency of carnot engine

"e=1-\frac{T_{sink}}{T_{source}}"

"e=1-\frac{273}{393}"


"e=\frac{120}{393}=0.305"

efficiency of carnot engine can be given by formula mention below

"e=\frac{W}{Q}" Where Q is heat rejected at sink

therefore "Q= \frac{W}{e}"


"Q=\frac{21000}{0.305}"


"Q=6405J"

This rejected heat will melt the ice and can be given by below

"Q={m}{L_{f}}"

where m is the mass of melted ice and Lf is latent heat of fusion.

therefore

"m=\frac{Q}{L_f}" value of latent heat of fusion for

ice Lf ="3.33\times10^5 J/Kg"


by putting the value of Q annd Lf


"m=\frac{6405}{3.33\times 10^5}"

"m=0.21328\space Kg"

"\colorbox{aqua}{m=0.21 Kg}"


therefore mass of melted ice is 0.21 Kg in 2 minutes.